# LV-UNet: A Lightweight and Vanilla Model for Medical Image Segmentation

### v1

@@ Abstract
Although the progress made by large models in computer vision, optimization challenges, the complexity of transformer models, computational limitations, and the requirements of practical applications call for simpler designs in model architecture for medical image segmentation, especially in mobile medical devices that require lightweight and deployable models with real-time performance. However, some of the current lightweight models exhibit poor robustness across different datasets, which hinders their broader adoption. This paper proposes a lightweight and vanilla model called LV-UNet, which effectively utilizes pre-trained MobileNetv3-Large models and introduces fusible modules. It can be trained using an improved deep training strategy and switched to deployment mode during inference, reducing both parameter count and computational load. Experiments are conducted on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvair-SEG datasets, achieving better performance compared to the state-of-the-art and classic models.
@@
Medical image segmentation, lightweight, vanilla, MobileNetv3-Large, deep training strategy
@@ I Introduction
Medical image segmentation, which aims to accurately delineate anatomical structures or abnormalities, is essential in disease diagnosis and treatment planning. Unlike traditional methods that relied on manual or semiautomatic methods, which were time-consuming, subjective, and prone to interobserver variability, computer-assisted techniques can be more efficient. With advances in deep learning, which utilize large annotated datasets and well-designed neural networks to learn complex image representations and automatically infer pixel-level labels, medical image segmentation has been revolutionized.
@@
Recently, medical imaging solutions have been transitioning from laboratory settings to bedside environments. Termed point-of-care imaging, this approach involves performing testing and analysis directly alongside the patient, with the aim of improving patient care. This shift expects medical AI models to be more real-time and lightweight. For example, task shifting for point-of-care ultrasound (POCUS) includes applications in obstetrics, gynecology, emergency medicine, infectious diseases, and cardiac, abdominal, and pulmonary conditions. At the primary care level in resource-limited settings, it has the potential to expand diagnostic imaging capacity and lead to meaningful health outcomes [1]. AI systems for the point-of-care require considerations for lightweight design, deployability, and real-time performance. For example, Dulmage et al. [2] developed a point-of-care and real-time AI system to support clinician diagnosis of a wide range of skin diseases. Lightweight and fast segmentation is crucial in these scenarios, as it ensures the AI system can quickly and accurately delineate areas of interest, facilitating efficient operation on resource-constrained devices and providing precise, timely results.
@@
UNet [3] is a classic medical image segmentation model consisting of an encoder and a decoder with skip-connections, showing a significant improvement compared to previous attempts. UNet’s variants like Attention-UNet [4], U-Net++ [5], ResUnet [6], Unet 3+[7] and so on have achieved great success. Due to the impressive performance of Transformers in visual tasks in recent years, Transformer-based U-Net variants like Trans-UNet[8], Swin-UNet [9], Unetr[10] and so on have gained wider recognition and have shown remarkable effectiveness. However, these models entail significant computational overhead and a large number of parameters, making them difficult to use in point-of-care applications. Thus, more and more effort is paid to designing lightweight models in medical image segmentation scenarios. Although existing lightweight medical image segmentation models have made significant progress, they often lack robustness across different datasets. How to address these issues by designing a model with reduced computational overhead, fewer parameters, faster inference time, and maintained performance, while also ensuring robustness across various datasets still remains a challenge.
@@
The model design of this paper is based on the following ideas: 1) Utilizing a pre-trained lightweight model in the encoder of a segmentation model is crucial, ensuring the model’s robustness across different datasets; 2) Combining pre-trained and expansion modules forms a straightforward approach to designing lightweight medical image segmentation models; 3) Introducing fusion modules enables further reduction of parameter count and computational load during the inference phase. We propose a UNet variant called LV-UNet with a lightweight and vanilla model design containing pre-trained weights from MobileNetv3-Large and fusible expansion modules, and also adopt the fusion method in VanillaNet and improve the deep training strategy. Experiments are conducted on the ISIC 2016 [11], BUSI [12], CVC-ClinicDB [13], CVC-ColonDB [14], and Kvair-SEG datasets [15], whose results show that the presented method has competitive performances compared to the state-of-the-art and classic methods.
@@ II-A Lightweight medical image segmentation models
Significant efforts have been devoted to designing lightweight models for medical image segmentation, with a primary focus on variants of U-Net. Valanarasu et al. [16] proposed UNeXt, a convolutional MLP-based network for image segmentation, featuring an initial convolutional stage and a latent MLP stage with tokenized MLP blocks that project convolutional features, reducing parameters and computational complexity while improving segmentation performance by shifting input channels to learn local dependencies. Ruan et al.[17] proposed a MAL-UNet that integrates several different modules: DGA for global and local feature extraction, IEA for dataset characterization and sample connectivity enhancement, CAB for multi-stage feature fusion generating channel-axis attention maps, and SAB for spatial-axis attention map generation on multi-stage features. Ruan et al. [18] proposed EGE-UNet, integrating Group multi-axis Hadamard Product Attention (GHPA) and Group Aggregation Bridge (GAB) modules to extract diverse pathological information and fuse multi-scale features effectively, significantly reducing the model size. Liu et al. [19] proposed Rolling-UNet, combining CNN with MLP to improve medical image segmentation by effectively capturing both local features and long-distance dependencies. Li et al. [20] introduced U-KAN, an enhanced U-Net variant integrating Kolmogorov-Arnold networks (KANs) [21] to improve accuracy and interpretability in medical image segmentation and diffusion models, which achieved superior performance with reduced computational cost compared to traditional U-Net approaches.
@@ II-B VanillaNet
”Vanilla” is a term often used metaphorically to describe something as plain, ordinary, or unadorned without anything redundant. Pursuing minimalism in design, VanillaNet [22] adopts a simple architecture that prioritizes computational efficiency and information preservation. By combining minimalistic operations, batch normalization, and activation functions, VanillaNet achieves competitive performance without excessive complexity. The deep training strategy and the series-informed activation function are keys to maintaining non-linearity in shallow networks.
@@ II-C MobileNetv3
MobileNetv3 is a state-of-the-art convolutional neural network designed for mobile and edge devices, combining high performance with efficiency. It builds on MobileNetv1 and MobileNetv2 [23] by integrating advanced techniques such as depthwise separable convolutions, inverted residuals with linear bottlenecks, and Squeeze-and-Excitation (SE) blocks. In addition, it introduces the H-Swish activation function, which improves computational efficiency and performance. MobileNetv3 uses neural architecture search (NAS) to optimize its structure for different performance and efficiency needs, resulting in two variants: MobileNetv3-Large for higher accuracy and MobileNetv3-Small for resource-constrained environments. It also incorporates training techniques like dropout, label smoothing, and mixup for better generalization.
@@ III-A The Role of Pre-training
Pre-training in large datasets allows models to learn rich and generalized feature representations, which significantly improves their performance in various tasks. These learned features capture important patterns and nuances in the data that may not be easily obtained through training on smaller datasets or from scratch. Lightweight models have fewer parameters, which can restrict their ability to learn and generalize from data effectively. Therefore, even though introducing complex changes to the architectures of model might seem beneficial, these modifications often do not provide enough improvement to outweigh the substantial performance boost gained from using pre-trained models. On the other hand, while complex model structures might offer improvements on specific datasets, such benefits are usually inconsistent. These intricate designs may overfit particular data characteristics, leading to poor generalization in different or unseen datasets. Relying on pre-trained weights is a more reliable approach to ensure the model’s overall performance and robustness.
III-B Efficient Model Deployment through Reparameterization
During the training phase, a more complex model can be utilized to ensure optimal learning and feature extraction. This model may include redundancy in parameters and use techniques like dropout and batch normalization for improved generalization. Once training is complete, the model undergoes reparameterization for the inference phase. This involves simplifying the model by compressing and fusing parameters, such as merging multiple convolutional filters into fewer ones. This reparameterized model maintains performance while significantly reducing computational and memory requirements, ensuring efficient deployment in resource-constrained environments.
@@ IV-A Overall Architecture
The model architecture contains an encoder, a decoder, and final output layers, which can be seen in Figure 1. The encoder includes pre-trained blocks from MobileNetv3-Large, as well as fusible encoder modules. The decoder consists of fusible decoder modules. There are skip connections between the corresponding layers in the encoder and decoder, and to avoid significantly increasing the number of parameters, element-wise addition is used instead of concatenation. The dimensions of each stage are [16, 24, 40, 80, 160, 240, 480].
Figure 1: The architecture of LV-UNet: the basic modules are pre-trained MobileNetv3-Large blocks(the initial convolution stage and the group \@slowromancapi@ to \@slowromancapiii@ (the first inverted residual block to ninth), fusible encoder blocks, fusible decoder blocks, skip-connections and the output block
@@ IV-B Pre-trained Modules in Encoder
LV-UNet uses MobileNetv3-Large from the initial convolutional stage to the ninth inverted residual block for the encoder block of LV-UNet. The initial convolutional stage receives the input image and performs basic feature extraction, and the others contain inverted residual blocks and depth-wise separable convolution layers, each with different convolution kernels and activation functions, progressively extracting more abstract features. We name fifteen inverted residual blocks in MobileNetv3-Large model as $r1$ to $r15$. These inverted residual blocks are grouped as follows: $r1$ and $r2$ as the first group -group \@slowromancapi@ , $r3$, $r4$, and $r5$ as the second group -group \@slowromancapii@, and $r6$, $r7$, $r8$, and $r9$ as the third group -group \@slowromancapiii@. We do not use the entire MobileNetv3-Large without fully connected layers, to allow flexible design of subsequent expansion modules. If we included inverted Residual block $r10$ through $r15$ after $r9$, the number of downsampling operations (and thus the image size constraints) would limit the fusion modules.
@@ IV-C1 Architecture
The training mode of a fusible encoder block contains a convolutional layer using a $1\times 1$ kernel with stride 1, a batch normalization operation, a Leaky ReLU function, a convolutional layer using a $1\times 1$ kernel, a pooling layer and a designed activation layer. The resolutions of output features of an encoder block are half of the input. In the deployment mode, the block only contains a convolutional layer, a pooling layer and the designed activation layer. A fusible decoder block is similar, where the pooling operation is replaced by an upsampling operation. The resolutions of output features of a decoder block are two times of the input. The architectures of the fusible blocks in the training and deployment modes are shown in Figure 2. The method to transfer the training mode to the deployment mode during the inference stage will be introduced in the section ”Deep Training Strategy and Deployment Mode”.
(a) The architecture of the fusible encoder block in the training and deployment modes
(b) The architecture of the fusible decoder block in the training and deployment modes
@@
IV-C2 Non-Linear Activation Layer
The Non-Linear Activation Function in LV-UNet follows the design in VanillaNet. Consider an input feature tensor, denoted as $x\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and number of channels, respectively. The activation function is formulated as follows:
@@ IV-C2 Non-Linear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. The Non-Linear Activation Function is followed by a convolution and a batch normalization operation. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and non-linear relationships within the input data.
IV-C3 Deep Training Strategy and Deployment Mode
@@
where $e$ is the current epoch and $E$ is the number of deep training epochs. At the beginning of the training, the activation function is a ReLU function, while at the end of the training, it is an identity mapping function. The nonlinearity gradually decreases. Unlike the design in [22] that defines $a$ as $a=\frac{e}{E}$, the designed method can make the transition from linear to non-linear show a trend of starting slow and accelerating, which aids in the convergence of training.
@@
By using the deep training strategy, LV-UNet can be switched to deployment mode in the inference stage, which reduces parameters and computational requirements a lot.
@@ V-A Datasets
UNet 0.8401 0.9116 0.6218 0.7575 0.7649 0.8638 0.7425 0.8387 0.7327 0.8440
Attention-UNet 0.8430 0.9137 0.6445 0.7755 0.7955 0.8839 0.7523 0.8487 0.7449 0.8512
UNet++ 0.8375 0.9097 0.6211 0.7568 0.7951 0.8832 0.7246 0.8294 0.7565 0.8598
UNext 0.8505 0.9181 0.6122 0.7437 0.7709 0.8660 0.6613 0.7774 0.6880 0.8115
MALUNet 0.8374 0.9098 0.5875 0.7318 0.4994 0.6565 0.5361 0.6860 0.5724 0.7219
EGEUNet 0.8380 0.9103 0.5571 0.6949 0.5190 0.6774 0.5744 0.7098 0.5342 0.6928
Rolling-UNet-M 0.8472 0.9160 0.6132 0.7434 0.7342 0.8435 0.5382 0.6773 0.6990 0.8211
U-KAN 0.8517 0.9192 0.6368 0.7668 0.8243 0.9017 0.7367 0.8353 0.7338 0.8438
MobileNetv3-Large-UNet 0.8602 0.9240 0.6226 0.7504 0.8347 0.9087 0.7802 0.8700 0.7804 0.8747

### v2

@@ Abstract
While large models have achieved significant progress in computer vision, challenges such as optimization complexity, the intricacy of transformer architectures, computational constraints, and practical application demands highlight the importance of simpler model designs in medical image segmentation. This need is particularly pronounced in mobile medical devices, which require lightweight, deployable models with real-time performance. However, existing lightweight models often suffer from poor robustness across datasets, limiting their widespread adoption. To address these challenges, this paper introduces LV-UNet, a lightweight and vanilla model that leverages pre-trained MobileNetv3-Large backbones and incorporates fusible modules. LV-UNet employs an enhanced deep training strategy and switches to a deployment mode during inference by re-parametrization, significantly reducing parameter count and computational overhead. Experimental results on ISIC 2016, BUSI, CVC-ClinicDB, CVC-ColonDB, and Kvair-SEG datasets demonstrate a better trade-off between performance and the computational load. The code will be released at https://github.com/juntaoJianggavin/LV-UNet.
@@
Medical image segmentation, lightweight, vanilla, MobileNetv3-Large, deep training strategy, re-parametrization
@@ I Introduction
Medical image segmentation, which aims to accurately delineate anatomical structures or abnormalities, is essential in disease diagnosis and treatment planning. Unlike traditional methods that relied on manual or semiautomatic methods, which were time-consuming, subjective, and prone to inter-observer variability, computer-assisted techniques can be more efficient. The advent of deep learning  [1], utilizing large annotated datasets and well-designed neural networks to infer pixel-level labels, has revolutionized the field.
@@
Recently, medical imaging solutions have been transitioning from laboratory settings to bedside environments. Termed point-of-care imaging, this approach involves performing testing and analysis directly alongside the patient, with the aim of improving patient care. This shift expects medical AI models to be more real-time and lightweight. For example, task shifting for point-of-care ultrasound (POCUS) includes applications in obstetrics, gynecology, emergency medicine, infectious diseases, and cardiac, abdominal, and pulmonary conditions. At the primary care level in resource-limited settings, it has the potential to expand diagnostic imaging capacity and lead to meaningful health outcomes [2]. AI systems for point-of-care require considerations for lightweight design, deployability, and real-time performance. For example, Dulmage et al. [3] developed a point-of-care and real-time AI system to support clinician diagnosis of a wide range of skin diseases. Lightweight and fast segmentation is crucial in these scenarios, as it ensures the AI system can quickly and accurately delineate areas of interest, facilitating efficient operation on resource-constrained devices and providing precise, timely results.
@@
UNet  [4] is a classic medical image segmentation model consisting of an encoder and a decoder with skip connections, showing a significant improvement compared to previous attempts. UNet’s variants like Attention-UNet  [5], U-Net++  [6], ResUnet  [7], Unet 3+ [8] and so on have achieved great success. Due to the impressive performance of Transformers in visual tasks in recent years, Transformer-based U-Net variants like Trans-UNet [9], Swin-UNet [10], MT-UNet [11] and Unetr [12] and so on have gained wider recognition and have shown remarkable effectiveness. However, these models entail significant computational overhead and a large number of parameters, making them difficult to use in point-of-care applications. Thus, more and more effort has been paid to designing lightweight models in medical image segmentation scenarios. Although existing lightweight medical image segmentation models have made significant progress, they often lack robustness across different datasets. How to address these issues by designing a model with reduced computational overhead, fewer parameters, faster inference time, and maintained performance, while also ensuring robustness across various datasets remains a challenge.
@@
The model design of this paper is based on the following ideas: 1) Utilizing a pre-trained lightweight model in the encoder of a segmentation model is crucial, ensuring the model’s robustness across different datasets, especially in small datasets; 2) combining pre-trained and expansion modules forms a straightforward approach to designing lightweight medical image segmentation models; 3) introducing fusible modules enables re-parametrization during the inference phase, which further reduces parameter count and computational load. Based on these, we propose a UNet variant called LV-UNet with a lightweight and vanilla model design containing pre-trained weights from MobileNetv3-Large and fusible expansion modules, and also adopt the fusion method in VanillaNet and improve the deep training strategy. Experiments are conducted on the ISIC 2016 [13], BUSI [14], CVC-ClinicDB [15], CVC-ColonDB [16], and Kvair-SEG datasets [17], whose results show that the presented method achieves a better trade-off between performance and the computational load.
@@ II-A Lightweight medical image segmentation models
Significant efforts have been devoted to designing lightweight models for medical image segmentation, with a primary focus on variants of U-Net. Valanarasu et al. [18] proposed UNeXt, a convolutional MLP-based network for image segmentation, featuring an initial convolutional stage and a latent MLP stage with tokenized MLP blocks that project convolutional features, reducing parameters and computational complexity while improving segmentation performance by shifting input channels to learn local dependencies. Ruan et al. [19] proposed a MAL-UNet that integrates several different modules: DGA for global and local feature extraction, IEA for dataset characterization and sample connectivity enhancement, CAB for multi-stage feature fusion generating channel-axis attention maps, and SAB for spatial-axis attention map generation on multi-stage features. Ruan et al. [20] proposed EGE-UNet, integrating Group multi-axis Hadamard Product Attention (GHPA) and Group Aggregation Bridge (GAB) modules to extract diverse pathological information and fuse multi-scale features effectively, significantly reducing the model size.
@@ II-B VanillaNet
Pursuing minimalism in design, VanillaNet  [21] adopts a simple architecture that prioritizes computational efficiency and information preservation. By combining minimalistic operations, batch normalization, and activation functions, VanillaNet achieves competitive performance without excessive complexity. The deep training strategy and the series-informed activation function are keys to maintaining non-linearity in shallow networks.
@@ II-C MobileNetv3
MobileNetv3  [22] is a state-of-the-art convolutional neural network designed for mobile and edge devices, combining high performance with efficiency. It builds on MobileNetv1 [23] and MobileNetv2 [24] by integrating advanced techniques such as depth-wise separable convolutions, inverted residuals with linear bottlenecks, and Squeeze-and-Excitation (SE) blocks. In addition, it introduces the H-Swish activation function, which improves computational efficiency and performance. MobileNetv3 uses neural architecture search (NAS) to optimize its structure for different performance and efficiency needs, resulting in two variants: MobileNetv3-Large for higher accuracy and MobileNetv3-Small for resource-constrained environments.
@@ III-A The Role of Pre-training
Pre-training in large datasets allows models to learn rich and generalized feature representations, which significantly improves their performance in various tasks. These learned features capture important patterns and nuances in the data that may not be easily obtained through training on smaller datasets or from scratch. Lightweight models have fewer parameters, which can restrict their ability to learn and generalize from data effectively. Therefore, even though introducing complex changes to model architectures might seem beneficial, these modifications often do not provide enough improvement to outweigh the substantial performance boost gained from using pre-trained models. On the other hand, while complex model structures might offer improvements on specific datasets, such benefits are usually inconsistent. Relying on pre-trained weights is a more reliable approach to ensure the model’s overall performance and robustness.
III-B Efficient Model Deployment through Re-parametrization
During the training phase, a more complex model can be utilized to ensure optimal learning and feature extraction. Once training is complete, the model undergoes re-parametrization for the inference phase. This involves simplifying the model by compressing and fusing parameters, such as merging multiple convolutional filters into fewer ones. This re-parameterized model maintains performance while significantly reducing computational and memory requirements, ensuring efficient deployment in resource-constrained environments.
@@ IV-A Overall Architecture
The model architecture of LV-UNet contains an encoder, a decoder, and final output layers, which can be seen in Figure 1. The encoder includes pre-trained blocks from MobileNetv3-Large, as well as fusible encoder modules. The decoder consists of fusible decoder modules. There are skip connections between the corresponding layers in the encoder and decoder, and to avoid significantly increasing the number of parameters, element-wise addition is used instead of concatenation. The dimensions of each stage are 16, 24, 40, 80, 160, 240 and 480.
Figure 1: The architecture of LV-UNet: the basic modules include pre-trained MobileNetv3-Large blocks(the initial convolution stage and the group \@slowromancapi@ to \@slowromancapiii@ (the first inverted residual block to ninth), fusible encoder blocks, fusible decoder blocks, skip-connections, and the output block.
@@ IV-B Pre-trained Modules in Encoder
LV-UNet uses MobileNetv3-Large from the initial convolutional stage to the ninth inverted residual block for the encoder block of LV-UNet. The initial convolutional stage receives the input image and performs basic feature extraction, and the following layers contain inverted residual blocks with depth-wise separable convolution layers, different convolution kernels and activation functions, progressively extracting more abstract features. We name inverted residual blocks in the MobileNetv3-Large model as $r1$ to $r15$. These inverted residual blocks are grouped as follows: $r1$ and $r2$ as the first group (group \@slowromancapi@) , $r3$, $r4$, and $r5$ as the second group (group \@slowromancapii@), and $r6$, $r7$, $r8$, and $r9$ as the third group (group \@slowromancapiii@). We do not use the entire feature extraction part of the MobileNetv3-Large model, to allow flexible design of subsequent expansion modules.
@@ IV-C1 Architecture
The training mode of a fusible encoder block contains a convolutional layer using a $1\times 1$ kernel with stride 1, a batch normalization operation, a Leaky ReLU function, a convolutional layer using a $1\times 1$ kernel, a pooling layer and a designed activation layer. The resolutions of output features of an encoder block are half of the input. In the deployment mode, the block only contains a convolutional layer, a pooling layer, and the designed activation layer. A fusible decoder block is similar, where the pooling operation is replaced by an upsampling operation. The resolutions of output features of a decoder block are two times of the input. The architectures of the fusible blocks in the training and deployment modes are shown in Figure 2. The method to transfer the training mode to the deployment mode during the inference stage will be introduced in the section ”Re-parametrization and Deployment Mode”.
(a) The architecture of the fusible encoder block in the training and deployment modes.
(b) The architecture of the fusible decoder block in the training and deployment modes.
@@
Figure 3: Example visualizations of segmentation results of different models.
IV-C2 Nonlinear Activation Layer
The nonlinear activation function in LV-UNet follows the design of series-informed activation functions in VanillaNet. Consider an input feature tensor, denoted as $x\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ represent the height, width, and number of channels, respectively. The activation function is formulated as follows:
@@ IV-C2 Nonlinear Activation Layer
where $A_{s}\left(x_{h,w,c}\right)$ denotes the output of the activation function at position $(h,w,c)$ in the input tensor $h$ ranges from $1$ to $H$, $w$ ranges from $1$ to $W$, and $c$ ranges from $1$ to $C$. The function aggregates inputs from neighboring positions $(i+h,j+w,c)$ within a range defined by $-n$ to $n$. The weights $a_{i,j,c}$ control the contribution of each input, and $A(\cdot)$ represents the activation function applied to each input, typically a nonlinear function like ReLU. Furthermore, $b_{c}$ represents a bias term specific to each channel $c$. The nonlinear activation function is followed by a convolution and a batch normalization operation. By utilizing this formulation, LV-UNet can capture and incorporate contextual information from neighboring positions, enabling the model to learn and represent complex patterns and nonlinear relationships within the input data.
IV-C3 Re-parametrization and Deployment Mode
@@
where $e$ is the current epoch and $E$ is the number of deep training epochs. At the beginning of the training, the activation function is a ReLU function, while at the end of the training, it is an identity mapping function. The nonlinearity gradually decreases. Unlike the design in  [21] that defines $a$ as $a=\frac{e}{E}$, the designed method can make the transition from linear to nonlinear show a trend of starting slow and accelerating, which aids in the convergence of training.
@@
By using the deep training strategy, LV-UNet can be switched to deployment mode in the inference stage, which reduces parameters and computational requirements.
@@ V-A Datasets
UNet [4] 0.8401 0.9116 0.6218 0.7575 0.7649 0.8638 0.7425 0.8387 0.7327 0.8440
Attention-UNet [5] 0.8430 0.9137 0.6445 0.7755 0.7955 0.8839 0.7523 0.8487 0.7449 0.8512
UNet++ [6] 0.8375 0.9097 0.6211 0.7568 0.7951 0.8832 0.7246 0.8294 0.7565 0.8598
UNeXt [18] 0.8505 0.9181 0.6122 0.7437 0.7709 0.8660 0.6613 0.7774 0.6880 0.8115
MALUNet [19] 0.8374 0.9098 0.5875 0.7318 0.4994 0.6565 0.5361 0.6860 0.5724 0.7219
EGEUNet [20] 0.8380 0.9103 0.5571 0.6949 0.5190 0.6774 0.5744 0.7098 0.5342 0.6928
U-KAN [25] 0.8517 0.9192 0.6368 0.7668 0.8243 0.9017 0.7367 0.8353 0.7338 0.8438
MobileNetv3-Large-UNet [22] 0.8602 0.9240 0.6226 0.7504 0.8347 0.9087 0.7802 0.8700 0.7804 0.8747